OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (7, 2) (5, 4)

OpenStudy (luigi0210):

Use the point-slope formula: \[\LARGE y-y_{1}=m(x-x_{1})\] m=slope: \[\LARGE \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]